### v1

@@ Appendix 0.B Hoeffding’s Inequality
Theorem 0.B.1 (Hoeffding’s Inequality)
@@
From the two Theorems 0.B.1 and 3.1 regarding the original and adaptive Hoeffding’s Inequality respectively, we may derive the following two corollaries.
@@
As Theorems 0.B.1 and 3.1, the estimation error $\varepsilon$ from the original and adaptive Hoeffding’s Inequality are $\sqrt{\frac{\log(2/\sigma)}{2n}}$ and $\sqrt{\frac{0.6\cdot\log(\log_{1.1}n+1)+1.8^{-1}\cdot\log\left(24/{\sigma}\right)}{n}}$, respectively. We may prove the former is smaller than the latter either analytically or empirically. The inequalities of Corollary 1 can be written:
@@
For both the original and adaptive Hoeffding’s Inequalities, the estimation errors (between the sample mean and population mean) $\varepsilon$ are monotonically decreasing to sample size $n$ and $\sigma$.
@@
Proof sketch: By taking the (partial) derivatives of the two analytical expressions of $\varepsilon$ r.w.t. $n$ and $\sigma$ respectively, we may establish negative results. ∎
Appendix 0.C Sequential Analysis Parameter Design
0.C.1 Design parameters and output of group sequential design
@@
0.C.2 Sample Size Calculation for a Continuous Endpoint
@@
0.C.3 Design Plan Parameters and Output for Means
@@
Appendix 0.D Experimental results
Figure 9: More results on ProTIP effectiveness (Part 1).
Figure 10: More results on ProTIP effectiveness (Part 2).

### v2

@@ Appendix 0.C Hoeffding’s Inequality
Theorem 0.C.1 (Hoeffding’s Inequality)
@@
From the two Theorems 0.C.1 and 3.1 regarding the original and adaptive Hoeffding’s Inequality respectively, we may derive the following two corollaries.
@@
As Theorems 0.C.1 and 3.1, the estimation error $\varepsilon$ from the original and adaptive Hoeffding’s Inequality are $\sqrt{\frac{\log(2/\sigma)}{2n}}$ and $\sqrt{\frac{0.6\cdot\log(\log_{1.1}n+1)+1.8^{-1}\cdot\log\left(24/{\sigma}\right)}{n}}$, respectively. We may prove the former is smaller than the latter either analytically or empirically. The inequalities of Corollary 1 can be written:
@@
For both the original and adaptive Hoeffding’s Inequalities, the estimation errors (between the sample mean and population mean) $\varepsilon$ are monotonically decreasing to sample size $n$ (for n >= 2) and $\sigma$.
@@
By taking the (partial) derivatives of the two analytical expressions of $\varepsilon$ r.w.t. $n$ and $\sigma$ respectively, we may establish negative results as what follows.
  1. 1.
    $\sqrt{\frac{\log(2/\sigma)}{2n}}$ partial derivative with respect to $n$:
    $\frac{\partial\varepsilon(n,\sigma)}{\partial n}$ $=\frac{\partial}{\partial n}\left(\sqrt{\frac{\log(2/\sigma)}{2n}}\right)$
    $=\frac{1}{2}\left(\frac{\log(2/\sigma)}{2n}\right)^{-\frac{1}{2}}\cdot\frac{\partial}{\partial n}\left(\frac{\log(2/\sigma)}{2n}\right)$
    $=\frac{1}{2\sqrt{\frac{\log(2/\sigma)}{2n}}}\cdot\frac{\partial}{\partial n}\left(\frac{\log(2/\sigma)}{2n}\right)$
    $=\frac{1}{2\sqrt{\frac{\log(2/\sigma)}{2n}}}\cdot\left(-\frac{\log(2/\sigma)}{2n^{2}}\right)$
    $=-\frac{\log(2/\sigma)}{4n^{2}\sqrt{\frac{\log(2/\sigma)}{2n}}}$
  2. 2.
    $\sqrt{\frac{\log(2/\sigma)}{2n}}$ partial derivative with respect to $\sigma$:
    $\frac{\partial\varepsilon(n,\sigma)}{\partial\sigma}$ $=\frac{\partial}{\partial\sigma}\left(\sqrt{\frac{\log(2/\sigma)}{2n}}\right)$
    $=\frac{1}{2}\left(\frac{\log(2/\sigma)}{2n}\right)^{-\frac{1}{2}}\cdot\frac{\partial}{\partial\sigma}\left(\frac{\log(2/\sigma)}{2n}\right)$
    $=\frac{1}{2\sqrt{\frac{\log(2/\sigma)}{2n}}}\cdot\frac{\partial}{\partial\sigma}\left(\frac{\log(2/\sigma)}{2n}\right)$
    $=\frac{1}{2\sqrt{\frac{\log(2/\sigma)}{2n}}}\cdot\left(-\frac{1}{2n\sigma\ln{10}}\right)$
    $=-\frac{1}{4n\sigma\ln{10}\sqrt{\frac{\log(2/\sigma)}{2n}}}$
  3. 3.
    $\sqrt{\frac{0.6\cdot\log(\log_{1.1}n+1)+1.8^{-1}\cdot\log\left(24/{\sigma}\right)}{n}}$ partial derivative with respect to $\sigma$:
    $\begin{split}\frac{\partial\varepsilon(n,\sigma)}{\partial\sigma}&=\frac{\partial}{\partial\sigma}\left(\sqrt{\frac{0.6\cdot\log(\log_{1.1}n+1)+1.8^{-1}\cdot\log\left(24/{\sigma}\right)}{n}}\right)\\ &=\frac{1}{2}\left(\frac{0.6\cdot\log(\log_{1.1}n+1)+1.8^{-1}\cdot\log\left(24/{\sigma}\right)}{n}\right)^{-\frac{1}{2}}\\ &\quad\times\frac{\partial}{\partial\sigma}\left(\frac{0.6\cdot\log(\log_{1.1}n+1)+1.8^{-1}\cdot\log\left(24/{\sigma}\right)}{n}\right)\\ &=\frac{1}{2\sqrt{\frac{0.6\cdot\log(\log_{1.1}n+1)+1.8^{-1}\cdot\log\left(24/{\sigma}\right)}{n}}}\\ &\quad\times\frac{\partial}{\partial\sigma}\left(\frac{0.6\cdot\log(\log_{1.1}n+1)+1.8^{-1}\cdot\log\left(24/{\sigma}\right)}{n}\right)\\ &=\frac{1}{2\sqrt{\frac{0.6\cdot\log(\log_{1.1}n+1)+1.8^{-1}\cdot\log\left(24/{\sigma}\right)}{n}}}\\ &\quad\times\left(-\frac{1}{1.8\sigma n\ln{10}}\right)\\ &=-\frac{1}{3.6\sigma n\ln{10}\sqrt{\frac{0.6\cdot\log(\log_{1.1}n+1)+1.8^{-1}\cdot\log\left(24/{\sigma}\right)}{n}}}\end{split}$
    Therefore, the partial derivatives with respect to $n$ and $\sigma$ are both less than zero.
  4. 4.
    $\sqrt{\frac{0.6\cdot\log(\log_{1.1}n+1)+1.8^{-1}\cdot\log\left(24/{\sigma}\right)}{n}}$ under the square root, we have $f_{1}(n)+f_{2}(n,\sigma)$:
    $f_{1}(n)=\frac{0.6\cdot\log(\log_{1.1}n+1)}{n}$
    $f_{2}(n,\sigma)=\frac{1.8^{-1}\cdot\log\left(\frac{24}{\sigma}\right)}{n}$
    $f_{2}(n,\sigma)$ partial derivative with respect to $n$:
    $\frac{\partial}{\partial n}f_{2}(n,\sigma)$ $=\frac{\partial}{\partial n}\left(\frac{1.8^{-1}\cdot\log\left(\frac{24}{\sigma}\right)}{n}\right)$
    $=\frac{-1.8^{-1}\cdot\log\left(\frac{24}{\sigma}\right)}{n^{2}}$
    Therefore, $f_{2}(n,\sigma)$ is monotonically decreasing.
    The derivative of the denominator of $f_{1}(n)$ is 1. For the numerator $f(n)=0.6\cdot\log(\log_{1.1}n+1)$,
    $f^{\prime}(n)$ $=\frac{d}{dn}\left(0.6\cdot\log(\log_{1.1}n+1)\right)$
    $=0.6\cdot\frac{1}{\log(1.1)\cdot(\log_{1.1}n+1)}\cdot\frac{1}{n}\cdot\frac{1}{\ln(10)}$
    $f^{\prime}(n)$ is monotonically decreasing. When $n\geq 2$, $f^{\prime}(n)\leq f^{\prime}(2)=0.165$. Therefore, the derivative of the numerator is smaller than that of the denominator, and $f_{1}(n)$ is monotonically decreasing.
Appendix 0.D Sequential Analysis Parameter Design
0.D.1 Design parameters and output of group sequential design
@@
0.D.2 Sample Size Calculation for a Continuous Endpoint
@@
0.D.3 Design Plan Parameters and Output for Means
@@
Appendix 0.E More Experimental Results
Figure 8: More results on ProTIP effectiveness for SD V1.5 Pert. Rate 10% (Part 1).
Figure 9: More results on ProTIP effectiveness for SD V1.5 Pert. Rate 10% (Part 2).
Figure 10: More results on ProTIP effectiveness for SD V1.4 Pert. Rate 10% (Part 1).
Figure 11: More results on ProTIP effectiveness for SD V1.4 Pert. Rate 10% (Part 2).
Figure 12: Results on ProTIP effectiveness for SDXL Turbo Pert. Rate 10% (Part 1).
Figure 13: Results on ProTIP effectiveness for SDXL Turbo Pert. Rate 10% (Part 2).